OpenStudy (anonymous):

Add (n^3 + 2n) + (2n6^3-n) a) n^3 + 4n b) 4n^3 c) 3n^3+ n d) 3n^3-n

jimthompson5910 (jim_thompson5910):

Combine like terms So n^3+2n^3 = 3n^3 and 2n+(-n) = 2n-n = n Which means that the final answer is 3n^3 + n